OpenStudy (anonymous):

Jason had 3/4 pound of salami. He used 1/3 pound for sandwiches.How many pounds of salami does he have left?

OpenStudy (snowfire):

Find the LCD and then subtract.

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4}-\cfrac{1}{3}=\cfrac{\square -\square }{{\bbox[2pt,white, border: 2px solid black]{ \text{LCD} }}}\)

OpenStudy (jdoe0001):

what would be the LCD?

OpenStudy (anonymous):

12?

OpenStudy (jdoe0001):

yeap... so it's 12... so \(\bf \cfrac{3}{4}-\cfrac{1}{3}=\cfrac{(12\div 4\cdot 3)-(12\div 3\cdot 1)}{12}\)

OpenStudy (anonymous):

??? im lost :/

OpenStudy (jdoe0001):

lost? what part?

OpenStudy (anonymous):

=(12÷4⋅3)−(12÷3⋅1) 12

OpenStudy (jdoe0001):

yeap... you divide the denominator of the fraction by the LCD, and the result of that, you multiply times the numerator of the same fraction

OpenStudy (jdoe0001):

thus, tis how you'd add/subtract any fractions

OpenStudy (jdoe0001):

\(\bf \cfrac{(12\div 4\cdot 3)-(12\div 3\cdot 1)}{12}\implies \cfrac{(3\cdot 3)-(4\cdot 1)}{12}\implies \cfrac{9-4}{12}\implies \cfrac{5}{12}\)

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

yw